A rectangle with 3 more than the width has a perimeter of 48?
Answer by checkley77(12844)

I WILL ASSUME THAT YOU MEAN THE LENGTH IS 3 MORE THAN THE WIDTH.
L=W+3
2L+2W=48
2(W+3)+2W=48
2W+6+2W=48
4W=48-6
4W=42
W=42/4
W=10.5 FOR THE WIDTH.
L=10.5+3=13.5 FOR THE LENGTH.
PROOF:
2*13.5+2*10.5=48
27+21=48
48=48